Circular No. 5922

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET KUSHIDA (1994a) The following additional precise positions have been reported:

1994 UT α_{2000} δ_{2000} Observer
Jan. 11.63889 9^{h}47^{m}24.^{s}17 +2^{o}43′01′′.5 Verveer
12.66042 9 47 50.88 +2 33 44.4 Smith
13.98907 9 48 21.03 +2 22 02.4 Casulli
13.99198 9 48 20.99 +2 22 00.4  _{”}

A. Verveer and T. Smith (Perth). 0.3-m astrograph. V. S. Casulli and S. Valentini (Colleverde di Guidonia).

Orbital calculations suggest that this is indeed a short-period comet, the following elements being from 29 observations, Jan. 9-13:

T = 1993 Dec. 12.993 TT  ω = 214.^{o}706
e = 0.63365  Ω = 245.781 }2000.0
q = 1.36606 AU  i = 4.167
a = 3.72881 AU n^{o} = 0.136883 P = 7.20 years
1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 8 9^{h}45.^{m}38 + 3^{o}19.^{′}8 0.499 1.399 139.^{o}2 27.^{o}4 10.9
18 9 49.39 + 1 53.0 0.499 1.429 147.3 21.9 11.0
28 9 49.36 + 1 13.0 0.511 1.467 156.1 15.8 11.2
Feb. 7 9 46.82 + 1 15.6 0.538 1.511 164.6 10.0 11.4
17 9 43.67 + 1 49.7 0.581 1.562 168.8  7.0 11.8
27 9 41.68 + 2 39.7 0.643 1.619 164.5  9.4 12.1

NOVA CASSIOPEIAE 1993 A. D. Scott and A. Evans, Keele University; and T. Geballe, Joint As-

tronomy Centre, write: “Spectra obtained on Jan. 10 with the U.K. Infrared Telescope (+ cooled grating spectrometer CGS4) in the K and L bands show that C I at 3.378 μm is comparable in strength with Paδ, as occurred in the dusty nova V842 Cen. The first overtone vibrational-rotational bands of ^{12}CO remain strong (IAUC 5916), and there is evidence in the L band for ‘PAH’-like features at 3.34, 3.4, and 3.46 μm; such features, also seen in V842 Cen (Wichmann et al. 1991, Infrared Spectral Regions of Stars), may arise from nucleation sites for carbon dust and suggest that the visual light curve of this nova will shortly enter a deep minimum.”

1994 January 14 Daniel W. E. Green